Fermi National Laboratory


Accelerator Update

Wednesday March 17
The day shift began with the third day of shutdown.

At 9:03 AM, CUB technicians reported that they switched the Booster 95 LCW system from heat exchanger #1 to exchanger #2.

A water tech reported at 1:40 PM that he would be turning off various Main Injector (MI) LCW pumps over the next two days to test configurations.

At 6:45 PM, the cryo system at the Tevatron (TeV) sector A4 warmed up, which allowed the vacuum work to begin.

Safety system techs finished safety system testing in Linac, Booster, and MI by 10 PM.

Thursday March 18
The midnight shift began with the fourth day of the shutdown.

At 8:07 AM, a tech began testing MI LCW pumps around the ring. This testing occurred all day.

Operators interlocked Linac at 10:30 AM.

Operations established beam to the Linac dump at 1:30 PM. A Linac expert began tuning the beamline.

Operators had the Booster interlocked by 5:53 PM and performed a safety system test. Then while trying to set up Booster to take beam, operators discovered sparking and smoke coming from a power supply (B:MV100). Operators called in an expert.

At 6:10 PM, operators turned off a Booster device (ORBMP) because its firing was too rapid. Operators contacted an expert who helped them configure a temporary fix at 6:47 PM in order to run beam.

A Power supply tech and an operator replaced a contactor for the Booster power supply MV100. Operations established beam to the Booster dump at 8 PM.

Friday March 19
The midnight shift began with the fifth day of shutdown. There was beam in Linac and Booster with Booster experts tuning.

At 1:04 AM, TeV experts reported that the A4 vacuum leak had returned. They said that two magnets, a dipole (A44-5) and a quadrupole (A45-1), would need replacing.

An expert repaired the Booster ORBMP module at 8:03 AM.

Plans for Today and this Weekend

Linac

  • On and running
Booster
  • All work should be completed by noon
  • An expert will conduct low intensity beam tuning this afternoon
  • High intensity tuning this evening
  • Aperture scans on Saturday (3/20/04)
Main Injector
  • Expect to startup on Saturday (3/20/04) evening
  • Sunday: establish 120 GeV beam, MI8 line tuning, P1 line tuning and optics, beam to the Antiproton Source (Pbar)

Pbar

  • Wants 120 GeV beam as soon as possible from MI
  • Conduct reverse proton studies
  • Check the motorized quadrupoles

TeV

  • The TeV hopes to turn on at 0000 hours on Thursday (3/25/04), conduct a startup check and then begin studies
  • First shot on Saturday (3/27/04)

Recycler

  • Hope to have shots of protons on Monday (3/22/04) and then conduct studies

Switchyard

  • Hope to have high intensity beam on Monday (3/22/04)

MiniBooNE

  • Working on isolating and repairing a cooling water leak

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/19/2004   email Fermilab